OpenStudy (anonymous):

z= -2 + 2i find z^24 using de moivre's theorem

OpenStudy (anonymous):

do you recall what \(-2+2i\) is in polar form?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

multiply the angle by 24, raise \(2\sqrt{2}\) to the power of 24, which is a ridiculously large number

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh thats easy thank you again!

OpenStudy (anonymous):

well you might want to change the angle

OpenStudy (anonymous):

change it to what?

OpenStudy (anonymous):

\[\frac{3\pi}{4}\times 24=18\pi\]

OpenStudy (anonymous):

and \(18\pi\) is conterminal with 0 so you might want to write \[(2\sqrt{2})^{24}\left(\cos(0)+i\sin(0)\right)\]

OpenStudy (anonymous):

okay